how many possible 5 card hands from a standard 52-card deck playing cards with one spade, two hearts and two diamonds
Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

how many possible 5 card hands from a standard 52-card deck playing cards with one spade, two hearts and two diamonds
----------------
1 spade = 13/52 = 1/4
1st heart = 13/51
2nd heart = 12/50
1st diamond = 13/49
2nd diamond = 12/48
---------------
Multiply those.

Answer by Edwin McCravy(20055)

how many possible 5 card hands from a standard 52-card deck
playing cards with one spade, two hearts and two diamonds,
Choose the 1 spade C(13,1) = 13 ways.
Choose the 2 hearts C(13,2) = 78 ways.
Choose the 2 diamonds C(13,2) = 78 ways.

(13)(78)(78) = 79092 ways.

Edwin